# Transforming Policy-Car Swerving for Mitigating Stop-and-Go Traffic Waves: A Practice-Oriented Jam-Absorption Driving Strategy

### v1

@@ VI Conclusion
Among these parameters, determining an appropriate JAD speed that does not trigger traffic breakdown remains challenging in practice. This issue requires further advances in traffic-flow stability research, particularly empirical studies. Nevertheless, this limitation does not prevent the practical implementation of the proposed JAD strategy. In practice, a relatively higher JAD speed can be adopted; based on our experience and as demonstrated in this paper, speeds such as 50 km/h or even 70 km/h (as exhibited by police-car swerving behavior in the real world shown in Figure 2) are likely to be feasible.
To ensure effectiveness and avoid unintended disturbances to the traffic system, it is essential to verify or predict, prior to dispatching a JAD vehicle, that the observed stop-and-go wave satisfies the required conditions. Specifically, the wave should be an isolated wave propagating downstream, rather than periodic oscillations associated with a fixed-location bottleneck. In short, a JAD vehicle should be dispatched only when a qualifying isolated stop-and-go wave is detected.
Overall, we believe that the present analysis substantially advances JAD studies from a purely theoretical concept toward practical application. In particular, we propose leveraging police-car swerving behavior to implement JAD, avoiding the common assumption in many previous studies that CAVs are required. While CAVs are often assumed in theoretical research nowadays, in reality, fully controllable vehicles that can participate in everyday traffic management are still not widely available. Therefore, with the proposed JAD strategy incorporating various practical considerations, the next step is real-world testing.

### v2

@@ VI Conclusion
Among these parameters, determining an appropriate JAD speed that does not trigger traffic breakdown remains challenging in practice. This issue requires further advances in traffic-flow stability research, particularly empirical studies. Nevertheless, this limitation does not prevent the practical implementation of the proposed JAD strategy. In practice, a relatively higher JAD speed can be adopted: based on our experience and as demonstrated in this paper, speeds such as 50 km/h or even 70 km/h (as exhibited by police-car swerving behavior in the real world shown in Figure 1) are likely to be feasible.
Learning lessons from SPECIALIST [10, 9], to ensure effectiveness and avoid unintended disturbances to the traffic system, it is essential to verify or predict, prior to dispatching a JAD vehicle, that the observed stop-and-go wave satisfies the required conditions. Specifically, the wave should be an isolated wave propagating downstream, rather than periodic oscillations associated with a fixed-location bottleneck. In short, a JAD vehicle should be dispatched only when a qualifying isolated stop-and-go wave is detected.
Overall, we believe that the present analysis substantially advances JAD studies from a purely theoretical concept toward practical application. In particular, we propose leveraging police-car swerving behavior to implement JAD, avoiding the common assumption in many previous studies that CAVs are required. While CAVs are often assumed in theoretical research nowadays, in reality, fully controllable vehicles that can participate in everyday traffic management are still rarely seen. With the proposed JAD strategy incorporating various practical considerations, the next step is real-world testing.
Acknowledgement
We would like to thank Dr. Andreas Hegyi at TU Delft for sharing his valuable insights and practical experience with Variable Speed Limits and SPECIALIST, which helped inspire the ideas behind this paper.